# AI Can Be Cognitively Biased: An Exploratory Study on Threshold Priming in LLM-Based Batch Relevance Assessment

### v1

@@ Abstract.
threshold priming, priming effect, cognitive bias, large language models, relevance assessment
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456ccs: Information systems Relevance assessmentccs: Computing methodologies Natural language processingccs: Applied computing Psychology
@@ 1. Introduction
In this study, inspired by previous work (e.g. Scholer et al., 2013; Eickhoff, 2018) in the Information Retrieval (IR) community, we investigate whether LLMs are influenced by the priming effect when making relevance judgments across multiple documents simultaneously. Given that the relevance labels obtained are often used for training IR algorithms and ranking IR systems, relevance labels influenced by cognitive biases could potentially lead to biased algorithms and suboptimal IR systems. Consequently, the IR community has conducted research on how cognitive biases of human assessors, such as the priming effect (Scholer et al., 2013), the anchoring effect (Shokouhi et al., 2015), the decoy effect (Eickhoff, 2018), affect relevance judgment outcomes. The priming effect is a psychological phenomenon where exposure to certain stimuli (such as words, images, or concepts) influences an individual’s subsequent behavior, judgments, and decisions unconsciously (Tulving and Schacter, 1990; Vanderwart, 1984; Herr, 1986). Figure 1 demonstrates an example of the priming effect. In the left-hand context, where the stimuli are all related to food, individuals are more likely to associate the stimulus with “soup.” Conversely, in the right-hand context, where the stimuli pertain to bathing, individuals are more likely to associate the stimulus with “soap.” Scholer et al. (Scholer et al., 2013) discovered that when assessors evaluating document relevance, the quality of previously assessed documents can serve as a threshold, triggering the priming effect: if the quality of earlier assessed documents is relatively low, assessors tend to assign higher scores to subsequent documents. As the IR community begins to use LLMs for automatic relevance annotation (Faggioli et al., 2023; Thomas et al., 2023; Upadhyay et al., 2024), it is necessary to investigate to what extent LLMs are influenced by various contextual factors when making relevance judgments, resulting in biased outcomes that reflect human cognitive biases.
To address this issue, in this study, we propose and seek to answer two research questions (RQs) with our experiments:
@@ 2.2. Cognitive Biases in Human Search Process
Humans need to seek information to support various decisions, ranging from everyday decisions such as choosing a good Chinese restaurant in Flushing to major decisions such as deciding whether to undergo a surgery to remove a tumor. In the process of information seeking and interacting with information retrieval systems, individual users also often engage in various decision-making processes, such as, whether to continue browsing the Search Engine Result Page (SERP) or to end the current search (Wicaksono and Moffat, 2018; Zhang et al., 2018; Chen et al., 2021), determining whether a document is useful in meeting their information needs (Mao et al., 2016, 2017), and evaluating the interaction experience with an search system (Jiang et al., 2015; Siro et al., 2023; Liu et al., 2018). Therefore, cognitive biases have been studied in a variety of works in the field of Information Retrieval (IR) (Azzopardi, 2021; Liu, 2023). For instance, how the recency effect and the reference-dependent effect influence users’ perceptions of search satisfaction (Liu and Han, 2020; Liu et al., 2019); and how the anchoring effect, decoy effect, expectation confirmation, and priming effect influence users’ judgments of document relevance (Shokouhi et al., 2015; Scholer et al., 2013; Eickhoff, 2018; Wang and Liu, 2023, 2024). Researchers have also explored the potential impact of cognitive biases on socio-political view (Knobloch-Westerwick et al., 2015; Kulshrestha et al., 2017) and health decisions (White and Horvitz, 2013; Zhang, 2012) during the information-seeking process. Additionally, they have investigated how to adjust oversimplified rational user models based on cognitive biases to design evaluation metrics that better characterize user judgments and interaction experience (Zhang et al., 2020; Chen et al., 2022, 2023, 2024; Liu and Shah, 2019). Among these studies, the work most closely related to our study is probably Scholer et al. (Scholer et al., 2013), arguing that assessors, when shown examples of low relevance items, tended to rate subsequent items more highly than if they were shown highly or moderately relevant items first.
@@ 3. Research Questions
To investigate to what extent are LLMs influenced by threshold priming when making batch relevance judgments of documents, in this study, we propose the following reseearch questions (RQs):
@@ 5. Experimental Result
Figure 4 shows the scores assigned to the documents in the epilogue by different models under different batch setups when the prologue is either LT or HT. From Figure 4 one can observe that regardless of the batch setup, the scores from GPT-3.5, GPT-4, and LLaMa2-13B are generally influenced by threshold priming. Specifically, when the relevance threshold of the documents in the prologue is higher, these models tend to assign lower scores to the documents in the epilogue, and this difference is statistically significant (p ¡ 0.001). This trend is consistent with the observations made by Scholer et al. (Scholer et al., 2013). Additionally, one can also observe that LLaMa2-13B exhibits a slighter difference in relevance assessment of documents in the epilogue under LT and HT compared to the GPT series.
@@ 6. Discussion
This work, inspired by insights from behavioral economics, investigated the threshold priming bias in LLMs within the context of information retrieval system evaluation. The findings of this work contribute to a deeper understanding of AI systems’ judgments and decision-making processes from an interdisciplinary perspective, and offer practical implications for designing fairer and human-centered unbiased AI systems.
@@
While our study has contributed valuable insights into the threshold priming effect in LLMs, it also presents certain limitations. One limitation is the restricted variety of topics and document batches tested, which may not fully capture the diversity of real-world applications. Moreover, the reliance on specific LLM models could skew the generalizability of our findings. However, these limitations do not substantially detract from the value of our experiments. To mitigate these issues, we have employed robust statistical methods to ensure the reliability of our findings and plan to extend our research to more diverse datasets, threshold priming and reference conditions, and additional LLM configurations in future studies. This approach will help us explore the broader applicability of our results, further refine our understanding of LLM behavior in complex decision-making scenarios, and better connect the knowledge about cognitive biases with AI bias mitigation techniques (Azzopardi and Liu, 2024).
@@ 7. Conclusion
This study, as an exploratory study, paves the path toward further explorations on threshold priming in LLMs in IR evaluation tasks. For instance, what are the boundaries for triggering threshold priming? How are LLM’s relevance and credibility labels influenced by threshold priming affect system rankings in offline evaluations? Additionally, as observed in our experiments, LLMs may be subject to cognitive biases beyond threshold priming, such as the anchoring bias, decoy effect, and reference dependence, during relevance assessment tasks. Future research directions include identifying and quantifying the biases (including the cognitively biased judgments) occurring in LLM’s decisions to better understand their limitations, and exploring methods to reduce these biases and facilitate more reliable human-AI collaboration across varying tasks and problem spaces, user cognitive states, and interaction modalities (Markwald et al., 2023; Liu et al., 2020; Kiesel et al., 2021).
Similar to searchers with cognitive limits (Liu, 2023), LLMs, contrary to the widely-accepted rationality assumptions, may also be ”boundedly rational” to some extent. Based on the results observed in our experiments, we suggest that the IR community have the LLM judge the same set of documents in different topics, random orders (i.e., multiple trials) as well as relevance levels from each trial. Additionally, future research should explore how prompting engineering and fine-tuning techniques can help identify and effectively mitigate the cognitive biases of LLMs.

### v2

@@ Abstract.
Threshold priming, priming effect, cognitive bias, large language models, relevance judgment, information retrieval evaluation
ccs: Information systems Relevance assessmentccs: Computing methodologies Natural language processingccs: Applied computing Psychology
@@ 1. Introduction
In this study, inspired by previous work (e.g. Scholer et al., 2013; Eickhoff, 2018) in the Information Retrieval (IR) community, we investigate whether LLMs are influenced by the priming effect when making relevance judgments across multiple documents simultaneously. Given that the relevance labels obtained are often used for training IR algorithms and ranking IR systems, relevance labels influenced by cognitive biases could potentially lead to biased algorithms and suboptimal IR systems. Consequently, the IR community has conducted research on how cognitive biases of human assessors, such as the priming effect (Scholer et al., 2013), the anchoring effect (Shokouhi et al., 2015), the decoy effect (Eickhoff, 2018), affect relevance judgment outcomes. The priming effect is a psychological phenomenon where exposure to certain stimuli (such as words, images, or concepts) influences an individual’s subsequent behavior, judgments, and decisions unconsciously (Tulving and Schacter, 1990; Vanderwart, 1984; Herr, 1986). Figure 1 demonstrates an example of the priming effect. In the left-hand context, where the stimuli are all related to food, individuals are more likely to associate the stimulus with “soup.” Conversely, in the right-hand context, where the stimuli pertain to bathing, individuals are more likely to associate the stimulus with “soap.” Scholer et al. (Scholer et al., 2013) discovered that when assessors evaluating document relevance, the quality of previously assessed documents can serve as a threshold, triggering the priming effect: if the quality of earlier assessed documents is relatively low, assessors tend to assign higher scores to subsequent documents. As the IR community begins to use LLMs for automatic relevance annotation (Faggioli et al., 2023; Thomas et al., 2023; Upadhyay et al., 2024), it is necessary to investigate to what extent LLMs are influenced by various contextual factors when making relevance judgments, resulting in biased outcomes that reflect human cognitive biases. Also, more broadly, it is essential to investigate the extent to which LLMs are ”rational” in automated judgment tasks, and if they make inaccurate judgments under human cognitive bias triggers. While LLMs could significantly reduce the cost of labeling for Machine Learning tasks both within and beyond IR, we also need to consider the efforts needed for bias mitigation and scalable AI audit.
To address the research gap above, we propose and seek to answer two research questions (RQs) with our experiments:
@@ 2.2. Cognitive Biases in Human Search Process
Humans need to seek information to support various decisions, ranging from everyday decisions such as choosing a good Chinese restaurant in Flushing to major decisions such as deciding whether to undergo a surgery to remove a tumor. In the process of information seeking and interacting with information retrieval systems, individual users also often engage in various decision-making processes, such as, whether to continue browsing the Search Engine Result Page (SERP) or to end the current search (Wicaksono and Moffat, 2018; Zhang et al., 2018; Chen et al., 2021; Liu et al., 2017), determining whether a document is useful in meeting their information needs (Mao et al., 2016, 2017), and evaluating the interaction experience with an search system (Jiang et al., 2015; Siro et al., 2023; Liu et al., 2018). Therefore, cognitive biases have been studied in a variety of works in the field of Information Retrieval (IR) (Azzopardi, 2021; Liu, 2023). For instance, how the recency effect and the reference-dependent effect influence users’ perceptions of search satisfaction (Liu and Han, 2020; Liu et al., 2019); and how the anchoring effect, decoy effect, expectation confirmation, and priming effect influence users’ judgments of document relevance (Shokouhi et al., 2015; Scholer et al., 2013; Eickhoff, 2018; Wang and Liu, 2023, 2024). Researchers have also explored the potential impact of cognitive biases on socio-political view (Knobloch-Westerwick et al., 2015; Kulshrestha et al., 2017), learning and interest development (Liu and Jung, 2021), and personal health decisions (White and Horvitz, 2013; Zhang, 2012) during the information-seeking process. Additionally, they have investigated how to adjust oversimplified rational user models based on cognitive biases to design evaluation metrics that better characterize user judgments and interaction experience (Zhang et al., 2020; Chen et al., 2022, 2023, 2024; Liu and Shah, 2019). Among these studies, the work most closely related to our study is probably Scholer et al. (Scholer et al., 2013), arguing that assessors, when shown examples of low relevance items, tended to rate subsequent items more highly than if they were shown highly or moderately relevant items first.
@@ 3. Research Questions
To investigate to what extent are LLMs influenced by threshold priming when making batch relevance judgments of documents, in this study, we propose the following research questions (RQs):
@@ 5. Experimental Result
Figure 4 shows the scores assigned to the documents in the epilogue by different models under different batch setups when the prologue is either low threshold (LT) or high threshold (HT). For each combination, we calculated the average relevance scores of two groups of epilogues under each theme, one processed by LT and the other by HT. We then conducted a dependent t-test (please note that both groups of epilogues were composed of the same documents in the same order).
From Figure 4 one can observe that regardless of the batch setup, the scores from GPT-3.5, GPT-4, and LLaMa2-13B are generally influenced by threshold priming. Specifically, when the relevance threshold of the documents in the prologue is higher, these models tend to assign lower scores to the documents in the epilogue, and this difference is statistically significant (p ¡ 0.001). This trend is consistent with the observations made by Scholer et al. (Scholer et al., 2013). Additionally, one can also observe that LLaMa2-13B exhibits a slighter difference in relevance assessment of documents in the epilogue under LT and HT compared to the GPT series.
@@ 6. Discussion
This work, inspired by insights from behavioral economics, investigated the threshold priming bias in LLMs within the context of information retrieval system evaluation. The findings of this work contribute to a deeper understanding of AI systems’ judgments and decision-making processes from an interdisciplinary perspective, and offer practical implications for designing effective bias mitigation strategies and human-centered unbiased AI systems.
@@
While our study has contributed valuable insights into the threshold priming effect in LLMs, it also presents certain limitations. One limitation is the restricted variety of topics and document batches tested, which may not fully capture the diversity of real-world applications. Moreover, the reliance on specific LLM models could skew the generalizability of our findings. However, these limitations do not substantially detract from the value of our experiments. To mitigate these issues, we have employed robust statistical methods to ensure the reliability of our findings and plan to extend our research to more diverse datasets, threshold priming and reference conditions, different prompt structures for assessment tasks, and additional LLM configurations in future studies. This approach will help us explore the broader applicability of our results, further refine our understanding of LLM behavior in complex decision-making scenarios, and better connect the knowledge about cognitive biases with AI bias mitigation techniques (Azzopardi and Liu, 2024; Liu and Azzopardi, 2024).
@@ 7. Conclusion
This study paves the path toward further explorations on threshold priming as well as other related biases and open questions of LLMs in IR evaluation tasks. For instance, what are the boundaries for triggering threshold priming? How are LLM’s relevance and credibility labels influenced by threshold priming affect system rankings in offline evaluations? Additionally, as observed in our experiments, LLMs may be subject to cognitive biases beyond threshold priming, such as the anchoring bias, decoy effect, and reference dependence, during relevance assessment tasks. Future research directions include identifying and quantifying the biases (including the cognitively biased judgments) occurring in LLM’s decisions to better understand their limitations, and exploring methods to reduce these biases and facilitate more reliable human-AI collaboration across varying tasks and problem spaces, user cognitive states, and interaction modalities (Markwald et al., 2023; Liu et al., 2020; Kiesel et al., 2021).
Similar to searchers with cognitive limits (Liu, 2023), LLMs, contrary to the widely-accepted rationality assumptions, may also be ”boundedly rational” to some extent. Based on the results observed in our experiments, we suggest that the IR community have the LLM judge the same set of documents in different topics, random orders (i.e., multiple trials) as well as relevance levels from each trial. Additionally, future research can explore how prompting engineering and fine-tuning techniques can help identify and effectively mitigate the cognitive biases of LLMs in automated judgment tasks, and how we can leverage the expertise of (instead of replacing) domain experts and information professionals in supervising and debiasing the judgments of LLM assessors in varying scenarios.